The equation of the best fit line relating x, the number of absences, to y, the final grade is ŷ=0.449x-30.27. Student A missed 7 classes while Student B missed 25 classes. Assuming that the scope of the model is between missing 0 classes to 30 classes, the predicted Final Grades for these grades are as follows
Answer #1

final grade : ŷ=0.449x-30.27

for A

ŷA = 0.449*7 - 30.27 = -27.127

predicted final grade for student A : ŷA = -27.127

for B

ŷB = 0.449*25 - 30.27 = -19.045

predicted final grade for student B : ŷB = -19.045

P.S. (please upvote if you find the answer satisfactory)